OpenStudy (anonymous):

Which number is between 0.6 and 7/9 ?

OpenStudy (anonymous):

A. 23/27 B. 20/27 C. 14/27 D. 1/3

OpenStudy (anonymous):

@Psymon

OpenStudy (shamil98):

0.6 = 6/10

OpenStudy (shamil98):

7/9 = 0.77777777

OpenStudy (shamil98):

20/27 = 0.74

OpenStudy (anonymous):

so which one is it

OpenStudy (anonymous):

@shamil98 its b right?

OpenStudy (psymon):

ALl of your possible answers can easily relate to 7/9 because of an easy common denominator conversion. 7/9, if we make it into a 27 denominator, is 21/27. That means A is gone immediately. 6/10 = 3/5, which is definitely over 1/3, meaning D is out. 14/27 is ever so slightly under 1/2, since 14/28 is one half. And since its under 1/2, its definitely under .6, leaving only B

OpenStudy (shamil98):

nin = luigi = shamil = psymon

OpenStudy (psymon):

= key lime pie

OpenStudy (anonymous):

aight can you guys help me on my other queston